OpenStudy (anonymous):

lim (1-cosx)^2 / x x-> 0

Parth (parthkohli):

\[{d \over dx}\left({(1 - \cos x)^2 \over x} \right) = \quad? \]

Parth (parthkohli):

Keep differentiating. :)

OpenStudy (anonymous):

umm could you please explain where you got the d/dx ?

Parth (parthkohli):

Sorry... wait

Parth (parthkohli):

\[\left({{d\over dx}(1 - \cos x)^2 \over {d \over dx}(x)}\right) \]

OpenStudy (anonymous):

@ParthKohli Please try to not use l'Hospital's for simple limits. Most people have just started calculus, currently. And, in addition, you need to differentiate both the numerator and denominator.

Parth (parthkohli):

Okay, then graphing?

Parth (parthkohli):

@LolWolf Yes... fixed that just before your reply :)

OpenStudy (anonymous):

im also fine with the law of limits but im not sure about the other one

OpenStudy (anonymous):

I'd recommend using the rules given (that can be proven geometrically). i.e. \[ \lim_{x\to0}\frac{1-\cos(x)}{x}=0 \]And \[ \lim_{x\to0}\frac{\sin(x)}{x}=1 \]

Parth (parthkohli):

Try graphing the function using technology, then see how the function goes towards 0.

hartnn (hartnn):

u can do it like this:  use the formula sin x/x for 1st fraction and directly substitute x=0 in 2nd fraction. did u get it?